OpenStudy (anonymous):

Solve for 3/4b-1/6=1/2 b. b=4/9 b=1/4 b=8/9 b=1/2

OpenStudy (anonymous):

First, subtract 3/4b from 1/2b, which would be 1/4b. Now you have 1/6 = 1/4b. What you would want to do is get rid of the variable by multiplying the whole fraction of 1/6 by 4. you get b = 2/3